D=F-(M+N) FIND F?
V=Vo-32t Find for t?
R=h(p+3q) Find for q?
Im stuck with these three. Can you help?

Answer by wuwei96815(245)

D=F-(M+N)
D=F-M-N
D+M+N=F

V=Vo-32t
V-Vo=-32t
V-Vo/-32=t

R=h(p+3q)
R/h=p+3q
R/h - p = 3q
(R/h - p)/3 = q

I hope those answers are right.